# MathVC: An LLM-Simulated Multi-Character Virtual Classroom for Mathematics Education

### v1

@@ 3.1 Experimental Setup
3.2 Experimental Results
@@ 3.3 Further Analysis and Discussion
Ablation Studies Evaluating Key Prompt Designs in MathVC We further conducted ablation studies to validate the key prompt designs in MathVC, hoping to share insights with future research that will simulate characters at the same complexity level. Our investigation reveals that (1) editing is superior to regenerating in the character schema modifier; (2) generating acts is preferable to solely selecting acts in the dialogue act generator; and (3) two-step generation is more effective than the direct generation in the response generator. The results and analysis are in Appendix E.
@@
Error Analysis Through analyzing the sampled dialogues, we identify two key issues of MathVC that can be addressed in the future. First, we observed that the LLM-simulated students sometimes cannot elaborate on their deliberation process. Rather, they tend to directly offer an opinion, which makes the conversation less engaging. Second, instructibility remains a major challenge in the character simulation. For example, even when a character has been simulated to take charge of a certain sub-task during the team organization, the LLM may still generate a response mistakenly representing the character’s actions in solving a different sub-task. We conjecture that this discrepancy is sourced from the LLM’s attention to dominantly the local context (e.g., the current sub-task focus), while instructions indicated in the relatively distance context are overlooked (e.g., Alice’s sub-task assignment at the beginning of the discussion). Future work could address these issues by designing an appropriate schema to encourage responses with details and dynamically store and retrieve information in both long- and short-term character memory.
@@
Preliminary Results with Humans in the Loop While we are not able to recruit middle-school students in our current research, we conducted a preliminary study showing the potential for real humans to engage in collaborative MM tasks with LLM characters. An example dialogue conducted by one author is shown in Appendix F. We found that the interaction between the human and the simulated student characters is natural and effective. It showcases the promise for future deployment with real middle-school students.
@@
Simulation with Varying Character Names, Genders, and Careers Our main evaluation has focused on the simulation of students at different MM skill levels, given that it is the most essential demand for MathVC. To further understand the impact of other traits, we present sampled dialogues with different character names (other than the “placeholders” Alice, Bob, and Charlie), genders (e.g., when all characters are females), and careers (e.g., when characters come from different grade levels or countries). Our ablation analysis as described in Appendix G shows that these characteristics have only a minor impact on the response, and we did not observe any ethical concerns (e.g., biases) in our examination. We include further discussions in the Ethics Statement.
Limitations and Future Work Like most existing LLM agents, MathVC faces the issues of token costs and latency, when it has to query an LLM multiple times for each character’s response generation. In our early study, we experimented with the relatively cheaper and faster GPT-3.5-turbo but encountered unsatisfactory results. On the other hand, the pipeline design of MathVC’s architecture could lead to error accumulation, i.e., when one module does not yield satisfying outcomes, the following ones suffer the same. However, overall MathVC still demonstrated a huge potential to be deployed to serve real students and be used as a tool for mathematics education. In the future, one important extension is to allow for “personalization” of MathVC, such as simulating student characters based on the specific human student’s need (e.g., setting up each character’s dialogue act frequency depending on what kinds of dialogue interaction can better scaffold the student’s problem modeling and solving). Finally, we clarify that the close-to-4 score of MathVC in Table 1 by no means implies a perfect performance with our system, and we believe that more effort is still required for the system to benefit student communities in practice. More systematic evaluation of both the system and the human student (e.g., whether they exhibit enhanced MM skills after interacting with MathVC) should be conducted in collaboration with experts from Mathematics Education.
@@ 4 Related Work
LLM Character Research There is a growing number of work employing LLMs to construct autonomous agents for playing specific roles (Wang et al., 2023b; Park et al., 2023), such as programmers (Qian et al., 2023; Hong et al., 2023). Some studies have found the characteristic alignment of LLMs is sometimes unsatisfactory (Fischer, 2023; Wang et al., 2024). Approaches to improving the characteristic alignment are mainly in two categories: fine-tuning LLMs (Shao et al., 2023; Zhou et al., 2023; Liu et al., 2023; Yu et al., 2024), which requires high-quality data, and carefully designing specific prompts and architectures (Yu et al., 2022; Li et al., 2023a; Tang et al., 2024; Hua et al., 2024). Given the challenge of gathering realistic data from middle-school students due to privacy concerns, we opted for the second line of approach. In particular, we propose a novel idea of leveraging symbolic representations to enhance characteristics alignment, and we explore an innovative application of simulating less knowledgeable populations (e.g., students with bad math skills) for mathematics education.
@@
LLMs in Mathematics Education Leveraging the mathematical reasoning ability of LLMs, numerous studies have explored how to enhance mathematics education with LLMs. McNichols et al. (2024) utilize LLMs to generate multiple-choice mathematical questions. Levonian et al. (2023) investigate how to select math problems to assist math tutoring. Bulusu et al. (2023) integrate LLMs and a math graphing tool to plot math figures by language. Macina et al. (2023) propose a framework for human teachers to interact with LLMs to collect data for math tutoring and the resulting data can be used to fine-tune smaller tutoring models. Most relevant to us are the works of Markel et al. (2023) and Lee et al. (2023), which construct platforms with LLM-simulated students for teaching assistants and teachers training. However, existing work has only focused on the vanilla simulation with humans-agents interaction, while we look into an innovative yet challenging setting of simulating multiple students in a collaborative problem-solving scenario with both human-agent and agent-agent interaction and carefully design the system to emulate authentic students. Our work will thus offer unique insights into the applications of LLMs to complicated multi-agent settings both in mathematics education and beyond.
5 Conclusion
@@
We present MathVC, a novel LLM-powered platform simulating middle-school students collaborating in mathematics problem-solving. To this end, we proposed the use of symbolic schema and meta planner to encourage simulation alignment. Our experimental results demonstrated the effectiveness of our simulation and the promise for MathVC to be deployed to enhance mathematics education in the future.
6 Ethics Statement
Our research is motivated by the pressing need to alleviate the dependence on teachers to orchestrate collaborative student learning for mathematics modeling (MM), a critical skill in STEM fields. To the best of our knowledge, MathVC is the very first LLM-powered multi-character virtual classroom designed for this purpose. We envision that such a platform can eventually be deployed as effective take-home exercises for students practicing their MM skills. This is particularly important for marginalized communities where only limited teachers and educational resources are available, and our system has the potential to become a tool to promote educational equity. In addition, we also identify a few other benefits of how MathVC could assist students in learning. For example, MathVC can reduce the pressure and anxiety of students for taking peer discussions, thus increasing their opportunities to participate. This is especially beneficial for students who may be reluctant to participate due to shyness, low self-esteem, language barriers, or poor academic performance. Our system thus promotes inclusion, allowing a greater variety of students to participate, thereby increasing diversity and breadth. Finally, we note that teachers have the flexibility to configure the student characters according to each human student’s background and performance, which can thus enable personalized learning experiences. In the future, one could also extend our system to automatically optimize the character configuration based on the student’s past performance.
Despite all the promises discussed above, however, we advocate careful discussion and comprehensive evaluation of any AI system’s safety before applying them to serve real humans. Due to this reason, in our project aiming for the first prototype system of a mathematics virtual classroom, we opted not to involve real middle-school students in the loop, before we can confirm the system’s safety. On the other hand, performing human subject studies with middle-school students requires collecting informed consent from their parents or guardians and obtaining permission from the school, which typically needs months of preparation time or longer. Therefore, our main evaluation was only conducted by examining the dialogues simulated among solely virtual characters.
As we discussed in Section 3.3, one extension of our system could be to enable richer characteristics simulation, including but not limited to gender, age, name, culture, language, and academic background. The significant advantage of such features is to allow for more engaging character simulation because, for example, students may enjoy the interaction with classmates sharing similar backgrounds with them. However, the introduction of these features may trigger biases in the LLM, such as leading to an unfair simulation of students from different backgrounds. When human students are exposed to such unfairly simulated characters, stereotypes may be exaggerated. In our analysis (Appendix G), we have not observed any such unethical simulation. However, future research following our work should remain cautious, ensuring that the simulated students objectively and fairly reflect the diversity presented in the real world.
7 Reproducibility Statement
We have documented the major implementation details, both in the main text and the appendix sections, including the input and output for all the modules and the prompts for key components (Appendix I). Our source code will be publicly available in the future. Our experimental setup is described in detail in the appendix. We will publicize our sampled data and LLM-generated dialogues with annotators’ scores and comments.
8 Acknowledgment
We were grateful to receive support from Microsoft’s Accelerating Foundation Models Research program (https://www.microsoft.com/en-us/research/collaboration/accelerating-foundation-models-research/) for conducting this research. We also appreciate computing resources from the Office of Research Computing (https://orc.gmu.edu) and funding support from the Department of Computer Science at GMU. W. Mifdal was sponsored by the Office of Student Creative Activities and Research (OSCAR) at GMU. We thank Dr. Anthony Eamonn Kelly (GMU, Educational Psychology) and Dr. Melissa A. Gallagher (University of Houston, Department of Curriculum & Instruction) for early discussions about the idea, and students in the GMU NLP group (https://nlp.cs.gmu.edu/) for their thoughtful comments.
Appendix A Workflow Algorithm of MathVC
The workflow algorithm of MathVC is shown in Algorithm 1.
Algorithm 1 MathVC Workflow
1:Characteristics $C$, Math Question $Q$, Answer $A$, Common Mistakes $M$, Dialogue Act Candidates $DAC$, End Stage $end$, $Max\ Conv.\ Turn$
2:Dialogue $D$
3:Initialize stage $\mathrm{s}\leftarrow 0$, $D\leftarrow[]$
4:$i\leftarrow 0$
5:while $i\leq Max\ Conv.\ Turn$ do
6:     if $i=0$ then
7:         $S_{t}\leftarrow\textsc{Task\ Schema\ Generator}(Q,A)$triangleright generate task schema
8:         $S_{c}\leftarrow\textsc{Character\ Schema\ Generator}(S_{t},M,C)$triangleright generate character schema
9:         $DA\leftarrow\textsc{Dialogue\ Act\ Generator}(S_{c},\mathrm{s},DAC[\mathrm{s}],D)$triangleright generate dialogue act
10:         $R\leftarrow\textsc{response\ generator}(S_{c},DA)$triangleright generate response
11:     else
12:         $Next\ Speaker\leftarrow\textsc{Next\ Speaker\ Predictor}(D)$triangleright get next speaker
13:         sDialogue Stage $\mathrm{s}\leftarrow\textsc{Dialogue\ Stage\ Monitor}(D,\mathrm{s})$triangleright stage update
14:         if $\mathrm{s}\ is\ end$ then
15:              break
16:         end if
17:         $S_{c}\leftarrow\textsc{Character\ Schema\ Modifier}(S_{c},D)$triangleright modity character schema
18:         $DA\leftarrow\textsc{Dialogue\ Act\ Generator}(S_{c},\mathrm{s},DAC[\mathrm{s}],D)$
19:         $R\leftarrow\textsc{Response\ Generator}(S_{c},DA,D)$
20:     end if
21:     $D\leftarrow D+[R]$
22:     if $Human\_Response$ then
23:         $D\leftarrow D+[Human\_Response]$
24:     end if
25:     $i\leftarrow i+1$
26:end while
Appendix B Mathematical Modeling Discussion Stages and Dialogue Acts
@@
Appendix C Example and Prompt for Character Schema Generation
@@
Appendix D Details of Experimental Setup
The GPT-4 we used was GPT-4-turbo version “11-06-preview”. The temperature is set to 0.7, top_p to 1, repetition_penalty to 1 and max_tokens to 1024. The student character implementation is based on a lightweight open-source platform Gentopia (Xu et al., 2023), on top of which we further developed the proposed meta planner to form MathVC. In our simulation, we have two different setups of simulated characters:
@@ Appendix D Details of Experimental Setup
Appendix E Ablation Studies Evaluating Key Prompt Designs in MathVC
@@
Appendix F Human Interaction
@@
Appendix G Characteristics Analysis
@@
From the example, although we changed the names, the responses are quite similar (except for negligible language variations). Overall, we observe that these characteristic variations other than the math skills only have a very minor impact on the LLM’s response. Most of the generations with different variations remain the same and none of the generations have been greatly changed. In particular, we did not observe any unfair simulation, such as simulating students with certain names or from a certain gender population with excessively good or bad behaviors. The reason could be that through reinforcement learning with human feedback (Ouyang et al., 2022), GPT-4 has been fine-tuned to avoid potential biases of these variations. Additionally, our prompts focus on discussing mathematical problems, without emphasizing backgrounds other than the MM skill levels of students. Under this well-defined scope, these background variations are not modeled as decisive factors in our current simulation.
@@
Appendix H Human Evaluation Guideline

### v2

@@ 3 System Evaluation
We conducted experiments using middle-school mathematics problems from the MAP project (MAP n.d.) and GSM8k (Cobbe et al. 2021). Questions from the MAP include both answers and common student mistakes, while those from the GSM8k only have answers; therefore, we manually annotate mistakes for them. Due to the cost of manual evaluation, we sample a set of questions. We used GPT-4o (version “gpt-4o-2024-05-13”) as the backbone LLM.
While we aim for a human subject study with real-life middle-school students to test our system, at the current stage, our primary goal is to showcase the feasibility of a virtual classroom. Therefore, our current evaluation focuses on whether the multiple (3 in our experiments) simulated LLM characters can reasonably collaborate on solving MM tasks during the evaluation. The evaluation is conducted by sampling multiple simulated dialogues among the simulated students and manually assessing their quality. In our experiment, 10 examples are tested. For each question on the test set, we consider two student character configurations with varying MM skill combinations.
3.1 Experimental Setup
Baseline Systems We compare MathVC with four baselines. (1) Vanilla Simulation follows the prior work (Park et al. 2023) in prompt engineering the profile instruction to an LLM. Specifically, we directly present the LLM the math question, the student characteristics, and the dialogue history, and then instruct the LLM to generate responses as a middle-school student. (2) Domain-Specified Simulation further augments the vanilla simulation with domain-specific knowledge in its prompt, including the stages and possible acts in each stage, common student mistakes, the language style of a middle-school student, and so on. To demonstrate the effectiveness of our character schema and the meta planner, we include two more variants on top of domain-specified simulation, i.e., (2.1) w/ only character schema, where we include our proposed character schema in the prompt to the LLM and allow for continual schema modification, and (2.2) w/ only meta planner, where we augment the domain-specified simulation with our meta planner. The planned stage is added to the prompt for character response generation.
Evaluation Metrics The human evaluation was conducted by two student annotators. Each of the two annotators labeled all the conversations. The annotators were instructed to rate each sampled dialogue on the degrees of characteristic alignment and conversational procedural alignment it reflected (1: worst; 4: best). We report an average score. Although the students probably have little experience with math discussions look like, we provided the annotators with detailed notes. We believe that genuine conversations should meet at least the two “alignment” goals we defined, i.e., consistency with each character’s properties and thoughts on the math problem (“characteristics alignment”) and showing smooth multiple-stage transition in the overall conversation (“conversational procedural alignment”). Therefore, we have provided detailed explanations only for these two points. The annotator has also scored based solely on these two aspects, and our paper reports only these two metrics. Other factors that may contribute to a conversation being realistic for a middle school student, such as tone, were not considered during annotation. The annotation guidelines are presented in Appendix G.
3.2 Evaluation Results
@@ 3.3 Further Analysis and Discussion
Ablation Studies Evaluating Key Prompt Designs in MathVC
@@
We further conducted ablation studies to validate the key prompt designs in MathVC, hoping to share insights with future research that will simulate characters at the same complexity level. Our investigation reveals that (1) editing is superior to regenerating in the character schema modifier; (2) generating acts is preferable to solely selecting acts in the dialogue act generator; and (3) two-step generation is more effective than the direct generation in the response generator. The results and analysis are in Appendix D.
Error Analysis
@@
Through analyzing the sampled dialogues, we identify two key issues of MathVC that can be addressed in the future. First, we observed that the LLM-simulated students sometimes cannot elaborate on their deliberation process. Rather, they tend to directly offer an opinion, which makes the conversation less engaging. Second, instructibility remains a major challenge in the character simulation. For example, even when a character has been simulated to take charge of a certain sub-task during the team organization, the LLM may still generate a response mistakenly representing the character’s actions in solving a different sub-task. We conjecture that this discrepancy is sourced from the LLM’s attention to dominantly the local context (e.g., the current sub-task focus), while instructions indicated in the relatively distance context are overlooked (e.g., Alice’s sub-task assignment at the beginning of the discussion). Future work could address these issues by designing an appropriate schema to encourage responses with details and dynamically store and retrieve information in both long- and short-term character memory.
Preliminary Results with Humans in the Loop
@@
While we are not able to recruit middle-school students in our current research, we conducted a preliminary study showing the potential for real humans to engage in collaborative MM tasks with LLM characters. An example dialogue conducted by one author is shown in Appendix E. We found that the interaction between the human and the simulated student characters is natural and effective. It showcases the promise for future deployment with real middle-school students.
Simulation with Varying Character Names, Genders, and Careers
@@
Our main evaluation has focused on the simulation of students at different MM skill levels, given that it is the most essential demand for MathVC. To further understand the impact of other traits, we present sampled dialogues with different character names (other than the “placeholders” Alice, Bob, and Charlie), genders (e.g., when all characters are females), and careers (e.g., when characters come from different grade levels or countries). Our ablation analysis as described in Appendix F shows that these characteristics have only a minor impact on the response, and we did not observe any ethical concerns (e.g., biases) in our examination. We include further discussions in the Ethics Statement.
@@ 4 Related Work
4.1 LLMs in Mathematics Education
@@
Leveraging the mathematical reasoning ability of LLMs, numerous studies have explored how to enhance mathematics education with LLMs. (McNichols et al. 2024) utilize LLMs to generate multiple-choice mathematical questions. (Levonian et al. 2023) investigate how to select math problems to assist math tutoring. (Bulusu et al. 2023) integrate LLMs and a math graphing tool to plot math figures by language. (Macina et al. 2023) propose a framework for human teachers to interact with LLMs to collect data for math tutoring and the resulting data can be used to fine-tune smaller tutoring models. Most relevant to us are the works of (Markel et al. 2023) and (Lee et al. 2023), which construct platforms with LLM-simulated students for teaching assistants and teachers training. However, existing work has only focused on the vanilla simulation with humans-agents interaction, while we look into an innovative yet challenging setting of simulating multiple students in a collaborative problem-solving scenario with both human-agent and agent-agent interaction and carefully design the system to emulate authentic students. Our work will thus offer unique insights into the applications of LLMs to complicated multi-agent settings both in mathematics education and beyond.
4.2 LLM Character Research
There is a growing number of work employing LLMs to construct autonomous agents for playing specific roles (Wang et al. 2023b; Park et al. 2023), such as programmers (Qian et al. 2023; Hong et al. 2023). Some studies have found the characteristic alignment of LLMs is sometimes unsatisfactory (Fischer 2023; Wang, Morgenstern, and Dickerson 2024). Approaches to improving the characteristic alignment are mainly in two categories: fine-tuning LLMs (Shao et al. 2023; Zhou et al. 2023; Liu et al. 2023; Yu et al. 2024), which requires high-quality data, and carefully designing specific prompts and architectures (Yu et al. 2022; Li et al. 2023; Tang et al. 2024; Hua et al. 2024). Given the challenge of gathering realistic data from middle-school students due to privacy concerns, we opted for the second line of approach. In particular, we propose a novel idea of leveraging symbolic representations to enhance characteristics alignment, and we explore an innovative application of simulating less knowledgeable populations (e.g., students with bad math skills) for mathematics education.
5 Conclusion and Discussion
@@
We present MathVC, a novel LLM-powered platform simulating middle-school students collaborating in mathematics problem-solving. To this end, we proposed the use of symbolic schema and meta planner to encourage simulation alignment. While LLM-based character simulations have been studied in prior work, the two alignment problems, i.e., characteristic alignment and conversational procedural alignment, have not been carefully discussed. Our experimental results demonstrated the effectiveness of our simulation and the promise for MathVC to be deployed to enhance mathematics education in the future. We believe our work will showcase the effective use of Generative AI for education and will inspire future research on the applications of multi-LLM agents to critical domains.
Like most existing LLM agents, MathVC faces the issues of token costs and latency, when it has to query an LLM multiple times for each character’s response generation. On the other hand, the pipeline design of MathVC’s architecture could lead to error accumulation. However, overall MathVC still demonstrated a huge potential to be deployed to serve real students and be used as a tool for mathematics education. In the future, one important extension is to allow for “personalization” of MathVC, such as simulating student characters based on the specific human student’s need (e.g., setting up each character’s dialogue act frequency depending on what kinds of dialogue interaction can better scaffold the student’s problem modeling and solving).
Finally, we clarify that the close-to-4 score of MathVC in Table 1 by no means implies a perfect performance with our system, and we believe that more effort is still required for the system to benefit student communities in practice. More systematic evaluation of both the system and the human student (e.g., whether they exhibit enhanced MM skills after interacting with MathVC) should be conducted in collaboration with experts from Mathematics Education.
6 Acknowledgement
The project is supported by Microsoft Research, NSF (#2418580, #2418582), and the Office of Research Computing at GMU.
Appendix A Mathematical Modeling Discussion Stages and Dialogue Acts
@@
Appendix B Example and Prompt for Character Schema Generation
@@
Appendix C Details of Experimental Setup
The GPT-4 we used was GPT-4-turbo version “11-06-preview”. The temperature is set to 0.7, top_p to 1, repetition_penalty to 1 and max_tokens to 1024. The student character implementation is based on a lightweight open-source platform Gentopia (Xu et al. 2023), on top of which we further developed the proposed meta planner to form MathVC. In our simulation, we have two different setups of simulated characters:
@@ Appendix C Details of Experimental Setup
Appendix D Ablation Studies Evaluating Key Prompt Designs in MathVC
@@
Appendix E Human Interaction
@@
Appendix F Characteristics Analysis
@@
From the example, although we changed the names, the responses are quite similar (except for negligible language variations). Overall, we observe that these characteristic variations other than the math skills only have a very minor impact on the LLM’s response. Most of the generations with different variations remain the same and none of the generations have been greatly changed. In particular, we did not observe any unfair simulation, such as simulating students with certain names or from a certain gender population with excessively good or bad behaviors. The reason could be that through reinforcement learning with human feedback (Ouyang et al. 2022), GPT-4 has been fine-tuned to avoid potential biases of these variations. Additionally, our prompts focus on discussing mathematical problems, without emphasizing backgrounds other than the MM skill levels of students. Under this well-defined scope, these background variations are not modeled as decisive factors in our current simulation.
@@
Appendix G Human Evaluation Guideline
@@
Appendix H Self-Play Annotation Details
Interrater agreement The average score for annotator 1 is 2.95 and for annotator 2 is 3.4, which indicates that annotator 1 is a relatively stricter evaluator. We also measure the agreement between them by seeing if they consistently agree on one system over the other (i.e., their ranks of systems). This is calculated with the average Spearman’s Rank Correlation coefficient (ranging from –1 to 1, with 1 indicating the same rank, 0 random, and –1 the completely opposite rank). We get 0.94 on the character alignment and 0.65 on the conversational alignment, which indicates an agreement of the ranking order between the two annotators.
Are the differences between the different versions of LLM agents evaluated statistically significant? We perform a t-test for the vanilla simulation, the domain-specified Simulation, and our method. The P-values are in the table 4 (P-value¡0.05 is statistically significant). We could see that except for the Conversational Procedural Alignment of the pair (vanilla simulation, domain-specific simulation), the other values show the. Specifically, our method is significantly better than the vanilla simulation in these two metrics.
Pair Characteristic Alignment P-value Conversational Procedural Alignment P-value
(Vanilla Simulation, Domain-Specified Simulation) 0.044 0.075
(Domain-Specified Simulation, Our method) 0.008 0.010
(Vanilla Simulation, Our method) <0.001 <0.001
Table 4: T-test